# On the Low-Temperature MCMC threshold:

### v1

@@ 9.2. Proof Strategy
We now explain the proof strategy we follow. For simplicity, we will use the notation of (3.3) in this proof section.
@@ 9.3.2. Two technical lemmas
Eq. 9.15 then follows. Combining Eqs. 9.14 and 9.15 with Eq. 9.13, we deduce that
@@ 10.1. Improving Previous Results
[Improved version of Theorem 2.1 of [GZ22a]] Suppose for some $\eta\in(0,1/3)$, $k=o(p^{1/3-\eta})$. Then:
@@
    Suppose further that $\sigma^{2}\leq 2k$. Then for every sufficiently large constant $D_{0}$, if $n\leq 2k\log(p/k)/(3\log D_{0})$, then w.h.p as $k$ increases, the cardinality of the set
    is at least $R^{n/5}$, where $R=\exp\left(\alpha k\log(p/k)/n\right)$ and $\alpha=(1+\eta)/(2+\eta)$.
@@ 10.2. Proof of the main negative result
In this section, we will apply Theorem 10.1 to prove Theorem 4.4. The crux of the argument can be described by decomposing the parameter space into a “low” overlap with $v^{*}$ part $A_{0}$, a “medium” overlap $A_{1}$ and a high overlap part $A_{1}.$ We then show that the probability mass placed on the low overlap set $A_{0}$ is much higher than that placed on the medium overlap set $A_{1}$. Since if we start from $A_{0}$ with a local Markov chain (such as the Metropolis chain on the Johnson graph) we have to go through the medium overlap set to get to the high overlap set $A_{2}$, this creates a bottleneck effect where the chain remains in the low overlap set for an exponential number of iterations.
The sets of low, medium, and high overlap with $v^{*}$ we fix throughout some arbitrary constants $0<\zeta_{1}<\zeta_{2}<1-\alpha$, where $\alpha$ is defined in Theorem 10.1, and then define,
@@
$\Gamma^{2}(\zeta^{*})\geq\zeta^{*}\exp\left(2(1-\zeta)\frac{k\log(p/k)}{n}\right)\Gamma^{2}(1)$
@@
$\frac{\pi_{\beta}(A_{1})}{\pi_{\beta}(A_{0})}\leq\exp\left(k\log\frac{pe}{k}-\frac{\alpha}{5}k\log\frac{p}{k}-\beta n(2k+\sigma^{2})M\right)$
where $M=\left[e^{-3}\zeta^{*}\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)-\exp\left(-\frac{2(1-\alpha)k\log(p/k)}{n}\right)\right]$
@@
$\pi_{\beta}(A_{1})$ $\leq\frac{\binom{p}{k}}{Z}\exp\left(-\beta\lVert Y-Xv\rVert^{2}_{2}\right)$ (10.6)
$\leq\frac{\binom{p}{k}}{Z}\exp\left(-\beta e^{-3}n\zeta^{*}(2k+\sigma^{2})\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)\right)$ (10.7)
@@
$\lVert Y-Xv\rVert^{2}<nR^{2}\Gamma^{2}(1)$
@@
$\pi_{\beta}(A_{0})$ $\geq\frac{R^{n/5}}{Z}\exp\left(-\beta nR^{2}\Gamma^{2}(1)\right)$ (10.10)
$=\frac{1}{Z}\exp\left(\frac{n}{5}\log R-\beta nR^{2}\Gamma^{2}(1)\right)$ (10.11)
@@
$\beta nR^{2}\Gamma^{2}(1)=\beta nR^{2}(2k+\sigma^{2})\exp\left(-\frac{2k\log(p/k)}{n}\right)$
@@
$\beta nR^{2}\Gamma^{2}(1)$ $=\beta n\exp\left(\frac{2\alpha k\log(p/k)}{n}\right)\left(2k+\sigma^{2}\right)\exp\left(-\frac{2k\log(p/k)}{n}\right)$
$=\beta n\left(2k+\sigma^{2}\right)\exp\left(-\frac{2(1-\alpha)k\log(p/k)}{n}\right)$
@@
$\pi_{\beta}(A_{0})\geq\frac{1}{Z}\exp\left(\frac{n}{5}\log R-\beta n\left(2k+\sigma^{2}\right)\exp\left(-\frac{2(1-\alpha)k\log(p/k)}{n}\right)\right)$
@@
$\pi_{\beta}(A_{0})\geq\frac{1}{Z}\exp\left(\frac{\alpha}{5}k\log(p/k)-\beta n\left(2k+\sigma^{2}\right)\exp\left(-\frac{2(1-\alpha)k\log p}{n}\right)\right)$
@@
$\frac{\pi_{\beta}(A_{1})}{\pi_{\beta}(A_{0})}\leq\exp\left(k\log\frac{pe}{k}-\frac{\alpha}{5}k\log\frac{p}{k}-\beta n(2k+\sigma^{2})M\right)$
for $M=\left[e^{-3}\zeta^{*}\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)-\exp\left(-\frac{2(1-\alpha)k\log(p/k)}{n}\right)\right]$
@@
$\beta>\frac{2}{n}\log(p/k)\exp\left(\frac{2k\log(p/k)}{n}\right)$
then with high probability,
@@
$\frac{\alpha}{5}k\log\frac{p}{k}+\beta n(2k+\sigma^{2})M-k\log\frac{pe}{k}$
for $M=\left[e^{-3}\zeta^{*}\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)-\exp\left(-\frac{2(1-\alpha)k\log(p/k)}{n}\right)\right]$. Since $\zeta^{*}<1-\alpha$ by assumption, the term $\exp\left(-\frac{2(1-\alpha)k\log(p/k)}{n}\right)$ in the square brackets is dominated by $e^{-3}\zeta^{*}\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)$. We also lower bound $2k+\sigma^{2}$ by $k$. It therefore suffices to show that
$\beta nk\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)-k\log\frac{pe}{k}=\Omega(k\log(p/k))$ (10.12)
@@
$\beta>\frac{2}{n}\log(p/k)\exp\left(\frac{2k\log(p/k)}{n}\right)$
@@
$\beta nk\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)-k\log\frac{pe}{k}>2k\log(p/k)-k\log\frac{pe}{k}=\Omega(k\log(p/k)$
@@ 11. Computational hardness of sparse tensor PCA
$\lambda_{ALG}(t,k)=\widetilde{\Theta}(\min\{k^{t/2},p^{t-1}/k^{t/2-1}\}).$
@@ 11.1. The hypergraph planted dense subgraph problem
$\rho_{\mathrm{ALG}}=\widetilde{\Theta}(\min\{1,p^{(t-1)/2}/k^{t-1}\}),$
@@ 11.2. Connecting back to sparse tensor PCA
Importantly to us, as long as $\rho=p^{-O(t)}$ holds, a simple average-case polynomial-time reduction can map any instance of HPDS with parameter $\rho$ to an instance of sparse tensor PCA with parameter $\lambda=\widetilde{\Theta}(k^{t/2}\log(1+2\rho))$ and $v=\mathbf{1}_{\mathcal{PC}}$ the indicator vector of $\mathcal{PC}.$ In particular, given this result, the computationally hard regime of HPDS $\rho=o(\rho_{\mathrm{ALG}})$ maps directly to the regime of sparse tensor PCA where $\lambda=o(\lambda_{ALG}(t,k))$ for some
$\lambda_{ALG}(t,k)=\widetilde{\Theta}(k^{t/2}\log(1+2\rho_{\mathrm{ALG}}))=\widetilde{\Theta}(k^{t/2}\rho_{\mathrm{ALG}})=\widetilde{\Theta}(\min\{k^{t/2},p^{t-1}/k^{t/2-1}\}),$
@@
The simple reduction works by employing the rejection kernel technique [BBH18]. Indeed by using the technique as described in [LZ22, Lemma 15] for probabilities 1/$21/2+\rho,1/$ 2 and some $\xi=\widetilde{\Theta}(\log(1+2\rho)/\sqrt{t})$, combined with the tensorization of the total variation distance [LZ22, Lemma 13], we directly get that the total variation between an instance $G$ from HPDS with parameter $\rho$ and an instance $Y$ from sparse tensor PCA for $\lambda=k^{t/2}\xi$ is $o(1)$, which completes the reduction.
@@ 12.2.1. The Pure Noise Model
Here there is no true vector $v^{*}$; we are simply trying to find the exactly $k$-sparse binary vector that minimises this squared error. Denote the optimal value of this optimization problem by $\psi_{2}$. Theorem 12.7 is our main result for this subsection. It improves Theorem A.1 of [GZ22b] by reducing the assumption that $\Omega(k\log k)=n=o(k\log p)$ (and hence $k=p^{o(1)}$) to $k=o(p^{1/3-\eta})$ for some $\eta>0$ and no assumption on $n$.
@@
$\binom{k}{\ell}\leq(\frac{ke}{\ell})^{\ell}$ (12.7)
@@
$\mathbb{E}\left(\min\left\{1,\Upsilon-1\right\}\right)$ $=\mathbb{E}_{Y}\left(\min\left\{1,\Upsilon-1\right\}\mathbf{1}\left(\min\left\{1-\Upsilon-1\right\}\geq\zeta^{n}\right)\right)$
$+\mathbb{E}_{Y}\left(\min\left\{1,\Upsilon-1\right\}\mathbf{1}\left(\min\left\{1-\Upsilon-1\right\}\leq\zeta^{n}\right)\right)$
@@
Since $\ell/e=\rho k/e\geq 5k/(6e)$ it therefore holds
@@
$-\frac{\log 4k}{n}+\left(\rho-(1-\alpha)\right)\frac{k\log\frac{p}{k}}{n}>\frac{1}{4}\frac{k\log\frac{p}{k^{2}}}{n}=\frac{1}{2}\log R_{1}$
@@
Next, we note that the $Y_{i}$’s are i.i.d., and $\frac{Y_{1}}{\sqrt{k}}\sim\mathcal{N}(0,\sigma^{2})$ and $1\leq\sigma^{2}\leq 3$. We apply Markov’s inequality to get
@@
$\mathbb{E}\left[\exp\left(\frac{Y_{1}^{2}}{12k}\right)\right]=\frac{1}{\sqrt{1-\frac{\sigma^{2}}{6}}}\leq\sqrt{2}$
@@
[New version of Lemma A.10 of [GZ22b]] Suppose that for some $\eta>0$, $4^{1+\eta}k^{3+\eta}\leq p$. Then, for all $\rho\in[\frac{1}{k},\rho_{*}]$ and $\zeta\in(0,1)$, it holds that for large enough $k$,
@@
Note that $\mathbb{E}\left(\frac{Y_{i}^{2}}{k}\right)=\sigma^{2}\in[1,3]$. From the moment generating function for a $\chi^{2}_{1}$ random variable, we have that
$\mathbb{E}\left[\exp\left(\frac{Y_{i}^{2}}{12k}\right)\right]=\left(1-\frac{\sigma^{2}}{6}\right)^{-\frac{1}{2}}\leq\sqrt{2}$
@@
By definition of $\rho_{*}$, we have $\rho_{*}=1-\log D/(6\log R_{1})$, so it holds that $1<\rho_{*}^{-1}\leq\frac{6}{5}$. We showed that for large enough $k$, $\log 4k/(\log(p/k^{2}))$ is at most $\frac{1}{2}$. We can therefore lower bound $f(\rho_{*})$ by
@@
Since $\zeta<1$, $\log\zeta<0$. It is therefore implied by $1<\rho_{*}^{-1}\leq\frac{6}{5}$ that we can lower bound $\rho_{*}\log\zeta$ by $2\log\zeta$:
@@
Since by assumption it holds that $k^{3+\eta}\leq p$, we have that for $\alpha=(1+\eta)/(2+\eta)<1$, it holds that
@@
then our goal is to obtain a lower bound on $Z_{t_{0}\sqrt{k}}$, where $Z_{t_{0}\sqrt{k}}$ denotes the cardinality of (12.15). Note that by (12.14), it holds that
$t_{0}\geq R^{1/2}\sqrt{1+\frac{\sigma^{2}}{k}}\exp\left(-\frac{k\log\frac{p}{k^{2}}}{2n}\right)\triangleq\tau$
so it suffices to lower bound $Z_{\tau\sqrt{k}}$. To see this, plug in $R=\exp\left(\alpha k\log(p/k)/n\right)$ to see that
$\tau$ $=\exp\left(\frac{\alpha k\log\frac{p}{k}}{2n}\right)\sqrt{1+\frac{\sigma^{2}}{k}}\exp\left(-\frac{k\log\frac{p}{k^{2}}}{2n}\right)$
$=\sqrt{1+\frac{\sigma^{2}}{k}}\exp\left(\frac{1}{n}\left[\frac{\alpha}{2}k\log\frac{p}{k}-\frac{1}{2}k\log\frac{p}{k^{2}}\right]\right)\leq t_{0}$
by (12.14). We let $D=D_{0}^{\frac{1}{2}}$ and then with this choice of $\tau$, all the conditions of Proposition 12.11 are satisfied ($\tau$ playing the role of $t$ in the proposition). We now require another improvement of a lemma from [GZ22b].
@@
$n^{-1}\log\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]\geq\frac{1}{4}\log R$
@@
$\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]=\binom{p}{k}\prod_{i=1}^{n}\mathbb{P}\left(\left|\frac{Y_{i}}{\sqrt{k}}-X\right|<\tau|Y\right)=\binom{p}{k}\prod_{i=1}^{n}p_{\tau,\frac{Y_{i}}{\sqrt{k}}}$
@@
$\log\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]=\log\binom{p}{k}+\sum_{i=1}^{n}\log p_{\tau,\frac{Y_{i}}{\sqrt{k}}}$

### v2

@@ 9.2. Proof Strategy
We now explain the proof strategy we follow. For simplicity, we will use the notation of (1.2) in this proof section, i.e. with
$Y=\lambda v^{\otimes t}/k^{t/2}+W$ (9.3)
where $W$ has i.i.d. $N(0,1)$ entries.
@@ 9.3.2. Two technical lemmas
Eq. 9.16 then follows. Combining Eqs. 9.15 and 9.16 with Eq. 9.14, we deduce that
@@ 10.1. Improving Previous Results
[Improved version of Theorem 2.1 of [GZ22a]] Suppose for some $\eta>0$, $k=o(p^{1/3-\eta})$. Then:
@@
    Suppose further that $\sigma^{2}\leq 2k$. Then for every sufficiently large constant $D_{0}$, if $n\leq\alpha k\log(p/k)/(3\log D_{0})$, then w.h.p as $k$ increases, the cardinality of the set
    is at least $R^{n/5}$, where $R=\exp\left(\alpha k\log(p/k)/n\right)$, $\alpha=(1+\eta)/(2+\eta)$, and $\mu=1-o(1)$ as $k\rightarrow\infty$.
@@ 10.2. Proof of the main negative result
In this section, we will apply Theorem 10.1 to prove Theorem 4.4. The crux of the argument can be described by decomposing the parameter space into a “low” overlap with $v^{*}$ part $A_{0}$, a “medium” overlap $A_{1}$ and a high overlap part $A_{2}$. We then show that the probability mass placed on the low overlap set $A_{0}$ is much higher than that placed on the medium overlap set $A_{1}$. Since if we start from $A_{0}$ with a local Markov chain (such as the Metropolis chain on the Johnson graph) we have to go through the medium overlap set to get to the high overlap set $A_{2}$, this creates a bottleneck effect where the chain remains in the low overlap set for an exponential number of iterations.
The sets of low, medium, and high overlap with $v^{*}$ we fix throughout some arbitrary constants $0<\zeta_{1}<\zeta_{2}<0.9-\alpha$, where $\alpha$ is defined in Theorem 10.1. Note $\zeta_{2}<\mu-\alpha-0.05$, since $\mu=1-o(1),$ as $k$ grows. Then define,
@@
$\Gamma^{2}(\zeta^{*})\geq\zeta^{*}\exp\left(2(1-\zeta^{*})\frac{k\log(p/k)}{n}\right)\Gamma^{2}(1)$
@@
$\frac{\pi_{\beta}(A_{1})}{\pi_{\beta}(A_{0})}\leq\exp\left(k\log\frac{pe}{k}-\frac{\alpha}{5}k\log\frac{p}{k}-\beta nM\right)$
where $M=\left[e^{-3}\zeta^{*}(2k+\sigma^{2})\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)-(2\mu k+\sigma^{2})\exp\left(-\frac{2(\mu-\alpha)k\log(p/k)}{n}\right)\right]$
@@
$\pi_{\beta}(A_{1})$ $\leq\frac{\binom{p}{k}}{Z}\exp\left(-\beta e^{-3}n\zeta^{*}(2k+\sigma^{2})\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)\right)$ (10.6)
@@
$\lVert Y-Xv\rVert^{2}<nR^{2}\Gamma^{2}(\mu)$
@@
$\pi_{\beta}(A_{0})$ $\geq\frac{R^{n/5}}{Z}\exp\left(-\beta nR^{2}\Gamma^{2}(\mu)\right)$ (10.9)
$=\frac{1}{Z}\exp\left(\frac{n}{5}\log R-\beta nR^{2}\Gamma^{2}(\mu)\right)$ (10.10)
@@
$\beta nR^{2}\Gamma^{2}(\mu)=\beta nR^{2}(2\mu k+\sigma^{2})\exp\left(-\frac{2\mu k\log(p/k)}{n}\right)$
@@
$\beta nR^{2}\Gamma^{2}(\mu)$ $=\beta n\exp\left(\frac{2\alpha k\log(p/k)}{n}\right)\left(2\mu k+\sigma^{2}\right)\exp\left(-\frac{2\mu k\log(p/k)}{n}\right)$
$=\beta n\left(2\mu k+\sigma^{2}\right)\exp\left(-\frac{2(\mu-\alpha)k\log(p/k)}{n}\right)$
@@
$\pi_{\beta}(A_{0})\geq\frac{1}{Z}\exp\left(\frac{n}{5}\log R-\beta n\left(2\mu k+\sigma^{2}\right)\exp\left(-\frac{2(\mu-\alpha)k\log(p/k)}{n}\right)\right)$
@@
$\pi_{\beta}(A_{0})\geq\frac{1}{Z}\exp\left(\frac{\alpha}{5}k\log(p/k)-\beta n\left(2\mu k+\sigma^{2}\right)\exp\left(-\frac{2(\mu-\alpha)k\log(p/k)}{n}\right)\right)$
@@
$\frac{\pi_{\beta}(A_{1})}{\pi_{\beta}(A_{0})}\leq\exp\left(k\log\frac{pe}{k}-\frac{\alpha}{5}k\log\frac{p}{k}-\beta nM\right)$
for $M=\left[e^{-3}\zeta^{*}(2k+\sigma^{2})\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)-(2\mu k+\sigma^{2})\exp\left(-\frac{2(\mu-\alpha)k\log(p/k)}{n}\right)\right]$
@@
$\beta>\frac{2}{c_{2}n}\log(p/k)\exp\left(\frac{2k\log(p/k)}{n}\right)$
for a small positive constant $c_{2}$, then with high probability,
@@
$\frac{\alpha}{5}k\log\frac{p}{k}+\beta nM-k\log\frac{pe}{k}$
for $M=\left[e^{-3}\zeta^{*}(2k+\sigma^{2})\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)-(2\mu k+\sigma^{2})\exp\left(-\frac{2(\mu-\alpha)k\log(p/k)}{n}\right)\right]$.
First, define $h$ as
$h=e^{-3}\zeta^{*}\left(2k+\sigma^{2}\right)-\frac{2\mu k+\sigma^{2}}{D_{0}^{\frac{6}{\alpha}(\mu-\alpha-\zeta^{*})}}$
Then we claim that $h$ is lower bounded by $c_{2}k$ for a small positive constant $c_{2}$ when $D_{0}$ is large enough. To see this, note that $e^{-3}\zeta^{*}\left(2k+\sigma^{2}\right)$ is lower bounded by $2e^{-3}\zeta^{*}k$, and we can upper bound the second summand of $h$ since $\sigma^{2}\leq 2k$ as follows,
$\frac{2\mu k+\sigma^{2}}{D_{0}^{\frac{6}{\alpha}(\mu-\alpha-\zeta^{*})}}\leq\frac{4}{D_{0}^{\frac{6}{\alpha}(\mu-\alpha-\zeta^{*})}}k$
Recall $\mu-\alpha-\zeta^{*}>0.05$ by construction, so for large enough $D_{0}$, $h$ is at least $c_{2}k$ for some small $c_{2}>0$.
Next, we claim that for large enough $D_{0}$, it holds that
$M\geq h\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)$
To prove this, observe that rearranging the definition of $h$, we get
$\frac{2\mu k+\sigma^{2}}{D_{0}^{\frac{6}{\alpha}(\mu-\alpha-\zeta^{*})}}=e^{-3}\zeta^{*}\left(2k+\sigma^{2}\right)-h$
and therefore,
$\frac{2\mu k+\sigma^{2}}{e^{-3}\zeta^{*}\left(2k+\sigma^{2}\right)-h}=D_{0}^{\frac{6}{\alpha}(\mu-\alpha-\zeta^{*})}$
Taking the logarithm of both sides, we see that
$\log\left[\frac{2\mu k+\sigma^{2}}{e^{-3}\zeta^{*}\left(2k+\sigma^{2}\right)-h}\right]=\frac{6}{\alpha}(\mu-\alpha-\zeta^{*})\log D_{0}$
and then by dividing across by $2(\mu-\alpha-\zeta^{*})$ and inverting both sides we get
$\frac{2(\mu-\alpha-\zeta^{*})}{\log\left[\frac{2\mu k+\sigma^{2}}{e^{-3}\zeta^{*}\left(2k+\sigma^{2}\right)-h}\right]}=\frac{\alpha}{3\log D_{0}}$
Since by assumption we have $n\leq\alpha k\log(p/k)/(3\log D_{0})$, it therefore holds that
$n\leq\frac{2(\mu-\alpha-\zeta^{*})k\log(p/k)}{\log\left[\frac{2\mu k+\sigma^{2}}{e^{-3}\zeta^{*}\left(2k+\sigma^{2}\right)-h}\right]}$
Rearranging, we get
$\log\left[\frac{e^{-3}\zeta^{*}\left(2k+\sigma^{2}\right)-h}{2\mu k+\sigma^{2}}\right]$ $\geq-\frac{2(\mu-\alpha-\zeta^{*})k\log(p/k)}{n}$
This gives us
$\log\left[e^{-3}\zeta^{*}\left(2k+\sigma^{2}\right)-h\right]-\frac{2\zeta^{*}k\log(p/k)}{n}$ $\geq\log\left[2\mu k+\sigma^{2}\right]-\frac{2(\mu-\alpha)k\log(p/k)}{n}$
Take the exponential of both sides to get
$\left[e^{-3}\zeta^{*}\left(2k+\sigma^{2}\right)-h\right]\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)$ $\geq\left[2\mu k+\sigma^{2}\right]\exp\left(-\frac{2(\mu-\alpha)k\log(p/k)}{n}\right)$
and finally, rearrange once again to get
$\left[e^{-3}\zeta^{*}\left(2k+\sigma^{2}\right)\right]\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)-\left[2\mu k+\sigma^{2}\right]\exp\left(-\frac{2(\mu-\alpha)k\log(p/k)}{n}\right)$
$\geq h\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)$
which proves the lower bound on $M$. Combining this with the fact that $h\geq c_{2}k$, it therefore suffices for Theorem 10.5 to show that
$\beta c_{2}nk\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)-k\log\frac{pe}{k}=\Omega(k\log(p/k))$ (10.11)
@@
$\beta>\frac{2}{c_{2}n}\log(p/k)\exp\left(\frac{2k\log(p/k)}{n}\right)$
@@
$\beta c_{2}nk\exp\left(-\frac{2\zeta^{*}k\log(p/k)}{n}\right)-k\log\frac{pe}{k}>2k\log(p/k)-k\log\frac{pe}{k}=\Omega(k\log(p/k)$
@@ 11. Computational hardness of sparse tensor PCA
$\lambda_{ALG}(t,k)=\tilde{\Theta}(\min\{k^{t/2},p^{(t-1)/2}/k^{t/2-1}\}).$
@@ 11.1. The hypergraph planted dense subgraph problem
$\rho_{\mathrm{ALG}}=\tilde{\Theta}(\min\{1,p^{(t-1)/2}/k^{t-1}\}),$
@@ 11.2. Connecting back to sparse tensor PCA
Importantly to us, as long as $\rho=p^{-O(t)}$ holds, a simple average-case polynomial-time reduction can map any instance of HPDS with parameter $\rho$ to an instance of sparse tensor PCA with parameter $\lambda=\tilde{\Theta}(k^{t/2}\log(1+2\rho))$ and $v=\mathbf{1}_{\mathcal{PC}}$ the indicator vector of $\mathcal{PC}.$ In particular, given this result, the computationally hard regime of HPDS $\rho=o(\rho_{\mathrm{ALG}})$ maps directly to the regime of sparse tensor PCA where $\lambda=o(\lambda_{ALG}(t,k))$ for some
$\lambda_{ALG}(t,k)=\tilde{\Theta}(k^{t/2}\log(1+2\rho_{\mathrm{ALG}}))=\tilde{\Theta}(k^{t/2}\rho_{\mathrm{ALG}})=\tilde{\Theta}(\min\{k^{t/2},p^{(t-1)/2}/k^{t/2-1}\}),$
@@
The simple reduction works by employing the rejection kernel technique [BBH18]. Indeed by using the technique as described in [LZ22, Lemma 15] for probabilities 1/$21/2+\rho,1/$ 2 and some $\xi=\tilde{\Theta}(\log(1+2\rho)/\sqrt{t})$, combined with the tensorization of the total variation distance [LZ22, Lemma 13], we directly get that the total variation between an instance $G$ from HPDS with parameter $\rho$ and an instance $Y$ from sparse tensor PCA for $\lambda=k^{t/2}\xi$ is $o(1)$, which completes the reduction.
@@ 12.2.1. The Pure Noise Model
Here there is no “true” vector $v^{*}$; we are simply trying to find the exactly $k$-sparse binary vector that minimises this squared error. Denote the optimal value of this optimization problem by $\psi_{2}$. Theorem 12.7 is our main result for this subsection. It improves Theorem A.1 of [GZ22b] by reducing the assumption that $\Omega(k\log k)=n=o(k\log p)$ (and hence $k=p^{o(1)}$) to $k=o(p^{1/3-\eta})$ for some $\eta>0$ and no assumption on $n$.
@@
$\binom{k}{\ell}\leq\left(\frac{ke}{\ell}\right)^{\ell}$ (12.7)
@@
$\mathbb{E}\left(\min\left\{1,\Upsilon-1\right\}\right)$ $=\mathbb{E}_{Y}\left(\min\left\{1,\Upsilon-1\right\}\mathbf{1}\left(\min\left\{1,\Upsilon-1\right\}\geq\zeta^{n}\right)\right)$
$+\mathbb{E}_{Y}\left(\min\left\{1,\Upsilon-1\right\}\mathbf{1}\left(\min\left\{1,\Upsilon-1\right\}\leq\zeta^{n}\right)\right)$
@@
Since $\ell/e=\rho k/e\geq 5k/(6e)$, it therefore holds
@@
$-\frac{\log(6/5)k+\log 4k}{n}+\left(\rho-(1-\alpha)\right)\frac{k\log\frac{p}{k}}{n}>\frac{1}{4}\frac{k\log\frac{p}{k^{2}}}{n}=\frac{1}{2}\log R_{1}$
@@
Next, we note that the $Y_{i}$’s are i.i.d., and $\frac{Y_{1}}{\sqrt{k}}\sim\mathcal{N}(0,\frac{\sigma^{2}}{k})$ and $1\leq\sigma^{2}/k\leq 3$. We apply Markov’s inequality to get
@@
$\mathbb{E}\left[\exp\left(\frac{Y_{1}^{2}}{12k}\right)\right]=\frac{1}{\sqrt{1-\frac{\sigma^{2}}{6k}}}\leq\sqrt{2}$
@@
[New version of Lemma A.10 of [GZ22b]] Suppose that for some $\eta>0$, $k=o(p^{1/3-\eta})$. Then, for all $\rho\in[\frac{1}{k},\rho_{*}]$ and $\zeta\in(0,1)$, it holds that for large enough $k$,
@@
Note that $\mathbb{E}\left(\frac{Y_{i}^{2}}{k}\right)=\frac{\sigma^{2}}{k}\in[1,3]$. From the moment generating function for a $\chi^{2}_{1}$ random variable, we have that
$\mathbb{E}\left[\exp\left(\frac{Y_{i}^{2}}{12k}\right)\right]=\left(1-\frac{\sigma^{2}}{6k}\right)^{-\frac{1}{2}}\leq\sqrt{2}$
@@
By definition of $\rho_{*}$, we have $\rho_{*}=1-\log D/(6\log R_{1})$, so it holds that $1<\rho_{*}^{-1}\leq\frac{6}{5}$. It holds that $\log 4k/(k\log(p/k^{2}))$ tends to zero as $k$ grows. We can therefore lower bound $f(\rho_{*})$ for large enough $k$ by
@@
Since $\zeta<1$, $\log\zeta<0$. It is therefore implied by $1<\rho_{*}^{-1}\leq\frac{6}{5}$ that we can lower bound $\rho_{*}^{-1}\log\zeta$ by $2\log\zeta$:
@@
Since by assumption it holds that for large enough $k$, $k^{3+\eta}\leq p$, we have that for $\alpha=(1+\eta)/(2+\eta)<1$, it also holds that
@@
then our goal is to obtain a lower bound on $Z_{t_{0}\sqrt{k}}$, where $Z_{t_{0}\sqrt{k}}$ denotes the cardinality of (12.15). We let $D=D_{0}^{\frac{1}{2}}$, and then all the conditions of Proposition 12.11 are satisfied. We now require another improvement of a lemma from [GZ22b].
@@
$n^{-1}\log\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]\geq\frac{1}{4}\log R$
@@
$\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]=\binom{p}{k}\prod_{i=1}^{n}\mathbb{P}\left(\left|\frac{Y_{i}}{\sqrt{k}}-X\right|<t_{0}|Y\right)=\binom{p}{k}\prod_{i=1}^{n}p_{t_{0},\frac{Y_{i}}{\sqrt{k}}}$
@@
$\log\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]=\log\binom{p}{k}+\sum_{i=1}^{n}\log p_{t_{0},\frac{Y_{i}}{\sqrt{k}}}$